Suppose the mean of the student completed the quiz is 30 minutes. Find the probability of a student completed the quiz in less than 20 minutes.


1
Expert's answer
2022-01-18T12:26:18-0500

μ = 30 , standard deviation is not given so let it be "\sigma=5"

"z=\dfrac{X-\mu}{\sigma}"

"P(X<20)=P\left(z<\frac{20-30}{5}\right)=P(z<-2)=0.02275"

Hence, the probability of a student completed the quiz in less than 20 minutes is 0.02275